A recent study of 300 patients found that of 100 alcoholic patients, 87 had evaluated cholesterol level, and of 200 nonacoholic patients, 43 had evaluated cholesterol levels, if a patient is selected at a random, find the probability that the patient is the following, a, an alcoholic with evaluated cholesterol level? b, a nonalcoholic? C,a nonalcoholic with non evaluated cholesterol level
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)

300 patients: 100A(87 with elevated CL), 200NA (43 with elevated CL, 157 w/o)
a patient is selected at a random from the 300 patients.:
P(A-elevated CL) = 87/300
P(NA) = 200/300 = 2/3
P(NA-non elevated CL) = 157/300

Answer by ikleyn(52781)

.

In your post,  replace the word  " evaluated "  by the word  " elevated ",  at each appearance.


Use the words correctly,  in accordance with their meaning.